# Multi-virtual braid groups

### v1

@@ 3.4. Symmetric multi-virtual pure braid group
We have
@@ 4. Some quotients and subgroups of multi-virtual braid group
The generators and defining relations for the pure braid groups $M_{k}WP_{n}$ and $M_{k}UP_{n}$, and for semi-pure $M_{k}WH_{n}$ and $M_{k}UH_{n}$ are stated as follows. The group $M_{k}WB_{n}$ is quotient of the group $M_{k}VB_{n}$ by relation $F1$. From the relation $\sigma_{i}\sigma_{i+1}\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\sigma_{i}\sigma_{i+1}$, we obtain the following relation:
@@
The group $M_{k}UP_{n}$ is quotient of group $M_{k}WP_{n}$ by relation $F2$. From the relation $\sigma_{i+1}\sigma_{i}\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\sigma_{i+1}\sigma_{i}$, we obtain the following relation:
@@
The group $M_{k}WB_{n}$ is quotient of the group $M_{k}VB_{n}$ by relation $F1$.
@@
The group $M_{k}UH_{n}$ is quotient of group $M_{k}WH_{n}$ by relation $F2$. From the relation $\sigma_{i+1}\sigma_{i}\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\sigma_{i+1}\sigma_{i}$, we obtain the following relations:
@@
5. Three strands multi-virtual braid group
To illustrate definitions and results of the previous sections, we consider 2-virtual 3-strand braid group, some its quotients and subgroups. As rule, 3-strand braid groups are different from the braid groups with bigger number of strands and are studied separately (see [6], [8], [12], [16]).
The group $MVB_{3}=M_{2}VB_{3}$ is generated by 6 elements:
$\sigma_{1},~{}~{}\sigma_{2},~{}~{}\rho_{1},~{}~{}\rho_{2},~{}~{}\tau_{1},~{}~{}\tau_{2},$
and is defined by relations
$\rho_{1}^{2}=\rho_{2}^{2}=\tau_{1}^{2}=\tau_{2}^{2}=e,$
Consider a homomorphism which is defined on the generators:
$\varphi_{3}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\varphi_{3}$ is the 2-virtual pure braid group $MVP_{3}$. Let us find its presentation. Using the Reidemeister-Schreier method, we see that this group is generated by elements
$\mu_{12}=\rho_{1}\tau_{1},~{}~{}\mu_{21}=\tau_{1}\rho_{1},~{}~{}\mu_{23}=\rho_{2}\tau_{2},~{}~{}\mu_{32}=\tau_{2}\rho_{2},$
$\mu_{13}=\rho_{2}(\rho_{1}\tau_{1})\rho_{2}=\rho_{2}\mu_{12}\rho_{2},~{}~{}\mu_{31}=\rho_{2}(\tau_{1}\rho_{1})\rho_{2}=\rho_{2}\mu_{21}\rho_{2}.$
From the relations
follows
$\lambda_{12}^{\rho_{2}\rho_{1}}=\lambda_{23},~{}~{}~{}\mu_{12}^{\rho_{2}\rho_{1}}=\mu_{23}$
and all other formulas of conjugations by elements $\rho_{i}$.
From relations $\tau_{1}^{2}=\tau_{2}^{2}=1$ follow relations
$\mu_{21}=\mu_{12}^{-1},~{}~{}\mu_{32}=\mu_{23}^{-1},~{}~{}\mu_{31}=\mu_{13}^{-1}.$
Hence, we can remove elements $\mu_{21}$, $\mu_{32}$ and $\mu_{31}$ from the list of generators $MVP_{3}$.
From relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$ using the rewritable process and formulas of conjugations by elements $\rho_{i}$ we get 6 relations:
$\lambda_{23}\lambda_{13}\lambda_{21}=\lambda_{21}\lambda_{13}\lambda_{23},~{}~{}~{}\lambda_{32}\lambda_{12}\lambda_{31}=\lambda_{31}\lambda_{12}\lambda_{32}.$
From relation $\sigma_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\sigma_{2}$ we do not have new relations.
From the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ follows relation
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
Hence, we have
Proposition 5.1.
The group $MVP_{3}$ is generated by elements
$\lambda_{12},~{}~{}\lambda_{21},~{}~{}\lambda_{13},~{}~{}\lambda_{31},~{}~{}\lambda_{23},~{}~{}\lambda_{32},~{}~{}\mu_{12},~{}~{}\mu_{13},~{}~{}\mu_{23}$
and is defined by the relations
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
5.1. 2-virtual semi-pure braid group on 3 strands
In this subsection we consider the map
$\psi_{3}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\psi_{3}$ is the 2-virtual semi-pure braid group $MVH_{3}$. Let us find its representation. This group is generated by elements
$x_{12}=\sigma_{1},~{}~{}x_{23}=\sigma_{2},~{}~{}x_{13}=\rho_{2}\sigma_{1}\rho_{2},$
$x_{21}=\rho_{1}\sigma_{1}\rho_{1},~{}~{}x_{32}=\rho_{2}\sigma_{2}\rho_{2},~{}~{}x_{31}=\rho_{2}\rho_{1}\sigma_{1}\rho_{1}\rho_{2},$
$z_{12}=\rho_{1}\tau_{1},~{}~{}z_{23}=\rho_{2}\tau_{2},~{}~{}z_{13}=\rho_{2}\rho_{1}\tau_{1}\rho_{2},$
$z_{21}=\tau_{1}\rho_{1},~{}~{}z_{32}=\tau_{2}\rho_{2},~{}~{}z_{31}=\rho_{2}\tau_{1}\rho_{1}\rho_{2}.$
Further, from relation $\tau_{i}^{2}=e$ follows relations $z_{21}z_{12}=z_{31}z_{13}=z_{32}z_{23}=e$.
From the relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$ follows 6 relation
$x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},$
$x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},$
$x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13}.$
From the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ follows relation
$z_{12}z_{13}z_{23}=z_{23}z_{13}z_{12}.$
The relations
$\sigma_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\sigma_{2},~{}~{}~{}\tau_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\tau_{2}$
give only conjugations rules by elements $\rho_{i}$.
Proposition 5.2.
  1. (1)
    The subgroup of $MVH_{3}$ which is generated by elements $x_{ij}$ is isomorphic to $VH_{3}$.
  2. (2)
    $\langle z_{12},z_{13},z_{23}~{}|~{}z_{12}z_{13}z_{23}=z_{23}z_{13}z_{12}\rangle\cong FVP_{3}$.
  3. (3)
    $MVH_{3}\cong VH_{3}*FVP_{3}$.
5.2. Some other maps
In this subsection we consider the map
$\chi_{3}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto e,~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
Denote its kernel $\ker\chi_{3}$ by $MVQ_{3}$ and find its presentation. It is easy to see that this group is generated by elements
$x_{12}=\sigma_{1},~{}~{}x_{23}=\sigma_{2},~{}~{}x_{13}=\rho_{2}\sigma_{1}\rho_{2},$
$x_{21}=\rho_{1}\sigma_{1}\rho_{1},~{}~{}x_{32}=\rho_{2}\sigma_{2}\rho_{2},~{}~{}x_{31}=\rho_{2}\rho_{1}\sigma_{1}\rho_{1}\rho_{2},$
$y_{12}=\tau_{1},~{}~{}y_{23}=\tau_{2},~{}~{}y_{13}=\rho_{2}\tau_{1}\rho_{2},$
$y_{21}=\rho_{1}\tau_{1}\rho_{1},~{}~{}y_{32}=\rho_{2}\tau_{2}\rho_{2},~{}~{}y_{31}=\rho_{2}\rho_{1}\tau_{1}\rho_{1}\rho_{2}.$
Further, from the relation $\tau_{i}^{2}=e$ follows the relations $y_{ij}^{2}=e$; from the relation $\sigma_{1}\sigma_{2}\sigma_{1}=\sigma_{2}\sigma_{1}\sigma_{2}$ follows 6 relations:
$x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},$
$x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},$
$x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13};$
from the relations $\rho_{1}\rho_{2}\sigma_{1}=\sigma_{2}\rho_{1}\rho_{2}$ and $\rho_{1}\rho_{2}\tau_{1}=\tau_{2}\rho_{1}\rho_{2}$ we do not have new relations; from the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ we have relations
$y_{12}y_{23}y_{12}=y_{23}y_{12}y_{23},~{}~{}y_{12}y_{13}y_{12}=y_{13}y_{12}y_{13},~{}~{}y_{13}y_{23}y_{13}=y_{23}y_{13}y_{23}.$
Proposition 5.3.
The group is $MVQ_{3}$ is generated by elements
$x_{12},~{}~{}x_{23},~{}~{}x_{13},~{}~{}x_{21},~{}~{}x_{32},~{}~{}x_{31},~{}~{}y_{12},~{}~{}y_{13}.~{}~{}y_{23},$
and is defined by relations
$(y_{12}y_{23})^{3}=(y_{12}y_{13})^{3}=(y_{23}y_{13})^{3}=e,$
$x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},$
$x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},$
$x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13}.$
Corollary 5.4.
$\langle x_{12},~{}~{}x_{23},~{}~{}x_{13},~{}~{}x_{21},~{}~{}x_{32},~{}~{}x_{31}\rangle\cong H_{1}*H_{2},$
where
$H_{1}=\langle x_{13},~{}~{}x_{21},~{}~{}x_{32}~{}|~{}x_{21}x_{13}x_{21}=x_{13}x_{21}x_{13},~{}~{}x_{13}x_{32}x_{13}=x_{32}x_{13}x_{32},~{}~{}~{}x_{32}x_{21}x_{32}=x_{13}x_{21}x_{13}\rangle,$
$H_{2}=\langle x_{12},~{}~{}x_{31},~{}~{}x_{23}~{}|~{}x_{12}x_{23}x_{12}=x_{23}x_{12}x_{23},~{}~{}x_{31}x_{12}x_{31}=x_{12}x_{31}x_{12},~{}~{}x_{23}x_{31}x_{23}=x_{31}x_{23}x_{31}\rangle.$
Corollary 5.5.
Let
$Y=\langle y_{12},~{}~{}y_{13},~{}~{}y_{23}~{}|~{}y_{12}^{2}=y_{13}^{2}=y_{23}^{2}=(y_{12}y_{13})^{3}=(y_{12}y_{23})^{3}=(y_{23}y_{13})^{3}=e\rangle.$
Hence, $MVQ_{3}\cong H_{1}*H_{2}*Y.$
Consider the map
$\chi_{4}\colon MVB_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\tau_{i}\mapsto e,~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\chi_{4}$ is denoted as $MVC_{3}$. It is easy to check
Proposition 5.6.
$MVC_{3}=VP_{3}*Y$.
5.3. Symmetric 2-virtual braid group on 3 strands
This group $\widetilde{MVB}_{3}$ is generated by 6 elements:
$\sigma_{1},~{}~{}\sigma_{2},~{}~{}\rho_{1},~{}~{}\rho_{2},~{}~{}\tau_{1},~{}~{}\tau_{2},$
and is defined by relations
$\rho_{1}^{2}=\rho_{2}^{2}=\tau_{1}^{2}=\tau_{2}^{2}=1,$
$\sigma_{1}\rho_{2}\rho_{1}=\rho_{1}\rho_{2}\sigma_{2},~{}~{}\tau_{1}\rho_{2}\rho_{1}=\rho_{2}\rho_{1}\tau_{2},$
and extra mixed relations
(13) $\sigma_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\sigma_{2},~{}~{}\tau_{1}\tau_{2}\rho_{1}=\rho_{2}\tau_{1}\tau_{2}.$
Let us defined a map on the generators
$\varphi_{3}\colon\widetilde{MVB}_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto\rho_{i},~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\varphi_{3}$ is the symmetric 2-virtual pure braid group $\widetilde{MVP}_{3}$. Let us find its presentation. Since it is the quotient of $MVP_{3}$ by the relations (13), we have rewrite these relations in the generators of $MVP_{3}$. From the relation $\sigma_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\sigma_{2}$ follows relation
$\lambda_{12}\,\mu_{13}\,\mu_{23}=\mu_{23}\,\mu_{13}\,\lambda_{12},~{}~{}\lambda_{21}\,\mu_{23}\,\mu_{13}=\mu_{13}\,\mu_{23}\,\lambda_{21},$
$\lambda_{23}\,\mu_{13}\,\mu_{12}=\mu_{12}\,\mu_{13}\,\lambda_{23},~{}~{}\lambda_{32}\,\mu_{12}\,\mu_{13}=\mu_{13}\,\mu_{12}\,\lambda_{32},$
$\mu_{12}\,\lambda_{13}\,\mu_{23}=\mu_{23}\,\lambda_{13}\,\mu_{12},~{}~{}\lambda_{31}\,\mu_{12}^{-1}\,\mu_{23}=\mu_{23}\,\mu_{12}^{-1}\,\lambda_{31}.$
From the relation $\tau_{1}\tau_{2}\rho_{1}=\rho_{2}\tau_{1}\tau_{2}$ follows that
$\langle\mu_{12},\mu_{13},\mu_{23}\rangle\cong\mathbb{Z}^{3}$
is a free abelian group of rank 3. From the relation $\tau_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\tau_{2}$ follows the relation
$\mu_{12}\mu_{13}\mu_{23}=\mu_{23}\mu_{13}\mu_{12}.$
5.4. Symmetric 2-virtual semi-pure braid group
Let
$\psi_{3}\colon\widetilde{MVB}_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto\rho_{i},~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2.$
The kernel $\ker\psi_{3}$ is the symmetric 2-virtual semi-pure braid group $\widetilde{MVH}_{3}$.
The group $\widetilde{MVH}_{3}$ is the quotient $MVH_{3}$ by the relations (13), we have rewritten these relations in the generators of $MVH_{3}$. From the relation $\sigma_{1}\tau_{2}\tau_{1}=\tau_{2}\tau_{1}\sigma_{2}$ follows the relations
$x_{21}z_{13}z_{23}=z_{13}z_{23}x_{21},~{}~{}x_{12}z_{23}z_{13}=z_{23}z_{13}x_{12},$
$x_{31}z_{12}z_{32}=z_{12}z_{32}x_{31},~{}~{}x_{32}z_{21}z_{31}=z_{21}z_{31}x_{32},$
$x_{13}z_{32}z_{12}=z_{32}z_{12}x_{13},~{}~{}x_{23}z_{31}z_{21}=z_{31}z_{21}x_{23}.$
From the relation $\tau_{1}\tau_{2}\rho_{1}=\rho_{2}\tau_{1}\tau_{2}$ follows that
$\langle z_{12},z_{13},z_{23}\rangle\cong\mathbb{Z}^{3}$
is a free abelian group of rank 3.
Remark 5.7.
If we consider a map
$\widetilde{MVB}_{3}\to S_{3},~{}~{}\sigma_{i}\mapsto e,~{}~{}\tau_{i}\mapsto e,~{}~{}\rho_{i}\mapsto\rho_{i},~{}~{}i=1,2,$
then from defining relations of $\widetilde{MVB}_{3}$ follows that $\rho_{1}=\rho_{2}$. Therefore this map does not form an epimorphism.
@@ Acknowledgments
This work has been supported by the grant the Russian Science Foundation, RSF 24-21-00102, https://rscf.ru/project/24-21-00102/.

### v2

@@ 3.4. Symmetric multi-virtual pure braid group
From the relation $\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(\beta)}=\rho_{i+1}^{(\beta)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}$, for $0<\alpha\neq\beta$, we obtain relation:
@@
From the relation $\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}\,\rho_{i}^{(0)}=\rho_{i+1}^{(0)}\,\rho_{i}^{(\alpha)}\,\rho_{i+1}^{(\alpha)}$, for $\alpha\neq 0$, we obtain relation:
$\lambda_{ij}^{(\alpha)}\lambda_{jk}^{(\alpha)}=\lambda_{jk}^{(\alpha)}\lambda_{ij}^{(\alpha)}.$
This last relation is already in $M_{k}VP_{n}$.
By the above calculations, we obtain
@@ 4. Some quotients and subgroups of multi-virtual braid group
The generators and defining relations for the pure braid groups $M_{k}WP_{n}$ and $M_{k}UP_{n}$, and for semi-pure $M_{k}WH_{n}$ and $M_{k}UH_{n}$ are stated as follows. The group $M_{k}WB_{n}$ is quotient of the group $M_{k}VB_{n}$ by relation $F1$.
From the relation $\sigma_{i}\sigma_{i+1}\rho_{i}^{(\alpha)}=\rho_{i+1}^{(\alpha)}\sigma_{i}\sigma_{i+1}$, we obtain the following relation:
@@
The group $M_{k}UP_{n}$ is the quotient of group $M_{k}WP_{n}$ by relation $F2$. From the relation $\sigma_{i+1}\sigma_{i}\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\sigma_{i+1}\sigma_{i}$, we obtain the following relation:
@@
The group $M_{k}WB_{n}$ is the quotient of the group $M_{k}VB_{n}$ by relation $F1$.
@@
The group $M_{k}UH_{n}$ is the quotient of group $M_{k}WH_{n}$ by relation $F2$. From the relation $\sigma_{i+1}\sigma_{i}\rho_{i+1}^{(\alpha)}=\rho_{i}^{(\alpha)}\sigma_{i+1}\sigma_{i}$, we obtain the following relations:
@@ Acknowledgments
The first and second authors are supported by the Ministry of Science and Higher Education of Russia (agreement No. 075-02-2026-1339). The third author gratefully acknowledges NBHM for the postdoctoral fellowship (Ref. No.: 0204/21(1)/2025-R& D-II/16279). The fourth author would like to acknowledge the support provided by the project CRG/2023/004921.
Conflict of Interest
The authors declare no conflicts of interest.